OpenStudy (anonymous):

Condense and/or solve: ln p+3 ln q+6 ln r

OpenStudy (anonymous):

My first thoughtful helper said it could not be condensed any further. Is this correct? If so perhaps the question is just asking me to solve? Can anyone solve thisS

OpenStudy (e.mccormick):

ln (p+3) ln (q+6) ln r or (ln p)+(3 ln q)+(6 ln r) ?

OpenStudy (anonymous):

the first one

OpenStudy (e.mccormick):

The they are multiplied and basic rules of logs can be used to combind them.

OpenStudy (e.mccormick):

Or.. let me see...

OpenStudy (agent0smith):

\[\Large \ln (p+3) \ln (q+6) \ln r\]If they're multiplied together, i don't see how to simplify. http://www.wolframalpha.com/input/?i=ln+%28p%2B3%29+ln+%28q%2B6%29+ln+r+simplify

OpenStudy (e.mccormick):

No... that would only be if they were multiplied INSIDE the ( ) Oops. Almost got that backwards myself. LOL

OpenStudy (e.mccormick):

Yah, ln(xy)=ln(x)+ln(y). So if it had been the second, there were a few things that could have been done. But for the first, no, that is it. And don't confuse the rules like I just did for a second. LOL. It happens, but try to make it as few times as possible.

OpenStudy (anonymous):

Thanks, lol I'm going to take a break but I will spend some more time with this one.

OpenStudy (e.mccormick):

OK, let me find a reference on the log rules. Might help for your review and understanding. http://www.purplemath.com/modules/logrules.htm